# GNSS-based Lunar Orbit and Clock Estimation With Stochastic Cloning UD Filter

### v1

@@ 2.4.2 Multi-Constellation Inter-System Biases (ISB)
The above measurement equations are expressed in ECEF coordinates and GPS Time (GPST). For Galileo, the broadcast clock correction is provided in Galileo System Time (GST) using the Galileo ephemeris parameters $\boldsymbol{\alpha}_{\mathrm{GAL}}$. We convert the Galileo transmitter clock bias to GPST via
@@ 2.4.3 Light-Time Equation
$\bar{t}_{tx,i}^{\mathrm{GPST}}=t_{k}^{\mathrm{GPST}}-\frac{\left\|\mathbf{r}_{k}^{\mathrm{ECEF}}-\bar{\mathbf{r}}_{tx,i}^{\mathrm{ECEF}}(\bar{t}_{tx,i}^{\mathrm{GPST}},\boldsymbol{\alpha})\right\|}{c},$ (20)
where $t_{k}^{\mathrm{GPST}}$ and $\mathbf{r}_{k}^{\mathrm{ECEF}}$ are obtained from the lunar time/frame quantities via the transformations in (10). The computation of $\bar{\mathbf{r}}_{tx,i}^{\mathrm{ECEF}}$ from broadcast ephemerides follows [IS-GPS-200N].
@@ 2.4.4 Ionospheric and Plasmaspheric Delays and Mitigation
GNSS signals propagating from Earth to lunar orbit experience dispersive delays through the ionosphere and plasmasphere. The dominant first-order term can be removed by forming an ionosphere-free (IF) combination of dual-frequency measurements. For example, using L1 and L5, the IF pseudorange is [kaplan2017understanding]
@@
where $\tilde{I}_{i,\mathrm{IF}}^{P}$ denotes the residual delay after first-order cancellation (e.g., higher-order dispersive terms and bending-related contributions), and
@@ 3 Stochastic-Cloning UD Filtering and Smoothing
This section presents a numerically robust filtering and smoothing framework for processing time-differenced carrier-phase (TDCP) measurements in low-observability lunar navigation scenarios. While stochastic-cloning Extended Kalman Filters (EKF) have previously been proposed to accommodate delayed-state measurements [Iiyama2024TDCPNavigation], we observe that their direct covariance-propagation formulations are susceptible to numerical instability, particularly when processing low-noise TDCP observables.
@@ 3.3.2 UD rank-one update
$\mathbf{f}=(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}_{k,m}^{\top},\qquad\mathbf{v}=\hat{\mathbf{D}}^{-}\mathbf{f}.$ (37)
@@
$\hat{\mathbf{U}}^{+}\hat{\mathbf{D}}^{+}\bigl(\hat{\mathbf{U}}^{+}\bigr)^{\top}$ $=\hat{\mathbf{U}}^{-}\left[\hat{\mathbf{D}}^{-}-\frac{1}{a}\mathbf{v}\mathbf{v}^{\top}\right]\bigl(\hat{\mathbf{U}}^{-}\bigr)^{\top}.$ (38)
@@ 1 $\mathbf{f}\leftarrow(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}^{\top}$;
2 $\mathbf{v}\leftarrow\hat{\mathbf{D}}^{-}\mathbf{f}$;
@@ 3.4 Reducing Computation Cost for the Measurement Update
We first process the $m_{1}$ delayed-state measurements. After these updates, we obtain UD factors $(\hat{\mathbf{U}}^{+},\hat{\mathbf{D}}^{+})$ for the augmented covariance as described in Section 3.3. We then process the remaining $m_{2}$ current-state-only measurements. For such measurements, $\mathbf{H}_{k,m}^{(2)}=\mathbf{0}$, so the auxiliary vectors in Algorithm 2,
$\mathbf{f}=(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}_{k,m}^{\top},\qquad\mathbf{v}=\hat{\mathbf{D}}^{-}\mathbf{f},$
inherit a block structure that does not require the bottom-right UD factor $\hat{\mathbf{U}}^{-}_{22}$:
$\mathbf{f}$ $=(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}_{k,m}^{\top}=\begin{bmatrix}(\hat{\mathbf{U}}^{-}_{11})^{\top}\,(\mathbf{H}^{(1)}_{k,m})^{\top}\\ (\hat{\mathbf{U}}^{-}_{12})^{\top}\,(\mathbf{H}^{(1)}_{k,m})^{\top}\end{bmatrix},\qquad\mathbf{v}=\hat{\mathbf{D}}^{-}\mathbf{f}=\begin{bmatrix}\hat{\mathbf{D}}^{-}_{11}(\hat{\mathbf{U}}^{-}_{11})^{\top}(\mathbf{H}^{(1)}_{k,m})^{\top}\\ \hat{\mathbf{D}}^{-}_{22}(\hat{\mathbf{U}}^{-}_{12})^{\top}(\mathbf{H}^{(1)}_{k,m})^{\top}\end{bmatrix}$, (46)
@@
Importantly, the covariance update equation (45) does not depend on $\hat{\mathbf{U}}^{+}_{22}$, and from (46), the matrix $\hat{\mathbf{U}}^{-}_{22}$ ($\hat{\mathbf{U}}^{+}_{22}$ from the previous measurement) is not involved in the computation of $\mathbf{f}$ and $\mathbf{v}$, which is used to compute $\hat{\mathbf{U}}^{+}_{11}$, $\hat{\mathbf{U}}^{+}_{12}$, $\hat{\mathbf{D}}^{+}_{11}$, and $\hat{\mathbf{D}}^{+}_{22}$ in Algorithm 2. In addition, the Kalman gain for the first half of the augmented state vector (i.e., the current state) can be computed without using the second half of the Kalman gain vector, which depends on $b_{j},(j=n+1,\ldots,2n)$ in Algorithm 2.
@@ 4.1.1 Time Update
$\frac{\partial f^{\text{dyn}}}{\partial\mathbf{x}}=F_{k}$ $=\begin{bmatrix}0_{3x3}&I_{3x3}&0&0&0&0_{3x1}\\ \frac{\partial\mathbf{a}}{\partial\mathbf{r}}&\frac{\partial\mathbf{a}}{\partial\mathbf{v}}&0&0&0&\frac{\partial\mathbf{a}}{\partial\gamma}\\ 0_{1x3}&0_{1x3}&0&1&0&0\\ \left(\frac{\partial(c\dot{\delta}t_{rel})}{\partial\mathbf{r}}\right)^{\top}&\left(\frac{\partial(c\dot{\delta}t_{rel})}{\partial\mathbf{v}}\right)^{\top}&0&0&0&0\\ 0_{1x3}&0_{1x3}&0&0&1&0\\ 0_{1x3}&0_{1x3}&0&0&0&0\\ 0_{1x3}&0_{1x3}&0&0&0&0\\ \end{bmatrix}$ (60)
@@
$Q_{pos,vel}=\begin{bmatrix}\frac{1}{3}q_{a}\Delta t^{3}I_{3x3}&\frac{1}{2}q_{a}\Delta t^{2}I_{3x3}\\ \frac{1}{2}q_{a}\Delta t^{2}I_{3x3}&q_{a}\Delta tI_{3x3}\end{bmatrix}$ (63)
@@ 4.1.2 Measurement Preprocessing
    Signal Acquisition and Tracking: The GNSS signals are acquired and tracked to extract raw pseudorange and carrier phase measurements. The detailed algorithms for weak GNSS signal acquisition and tracking are beyond the scope of this paper. Readers are referred to [Capuano2016StandaloneGPS, Simone2023LuGREReceiver] for more information.
@@
    where $R_{E}$ is the Earth’s mean radius, $\mathbf{r}_{tx}$ is the position of the transmitting GNSS satellite in ECEF coordinates, and $\mathbf{r}$ is the position of the lunar satellite receiver in ECEF coordinates. Remove all the measurements with tangential altitudes below a certain threshold mitigate the impact of ionospheric delays.
@@
    Compute the receiver noise of the pseudorange and carrier phase measurements based on the observed $C/N0$ values.
@@
    Compute the predicted measurements $\hat{z}_{k,i}$ and their residuals $y_{k,i}$ using the current state estimate $\hat{x}_{k|k-1}$:
@@
    This criterion discards the TDCP measurements whenever the residual exceeds a 3-sigma bound, or exceeds the wavelength of the signal $\lambda_{L}$ (=19 cm for L1 and 24 cm for L5), or the 3-sigma bound itself exceeds the wavelength, which indicates insufficient state knowledge and/or measurement noise to reliably detect cycle slips. For this work, we only discard the TDCP measurements with detected cycle slips, and do not attempt to repair them.
@@ 4.1.3 Measurement Update
$\mathbf{H}_{k,i}$ $=\begin{bmatrix}-\frac{(\mathbf{r}_{k}-\mathbf{r}_{tx,k,i})^{\top}}{\|\mathbf{r}_{k}-\mathbf{r}_{tx,k,i}\|}&0_{1x3}&1&0&0&0\end{bmatrix}\in\mathbb{R}^{1\times 10}$ (69)
@@ 5.1 Lunar Orbit
We used the initial conditions provided by [LCRNS2025] to simulate LCRNS satellites (LDN-1). Initial epoch of the simulation was set to March 1, 2025, 12:00:00 UTC, since the precise orbits, ephemeris, and the parameters needed for the ionospheric and plasmaspheric delay simulations were not available for the date in the reference (March 1, 2027, 00:00:00 UTC). To convert the initial conditions, the initial position and velocity vectors of the LCRNS satellite given at [LCRNS2025] were first converted to orbital elements at OP-frame (Orbital Plane frame) defined in [Ely2005FrozenLunarOrbits], and then the same initial orbital elements were used to generate the position and velocity vectors at the new initial epoch (March 1, 2025, 12:00:00 UTC) in the OP-frame. Finally, the position and velocity vectors were transformed back to the inertial frame (Moon-centered, axis aligned with the J2000 frame) for the orbit propagation. The initial orbital elements in the OP-frame are summarized in Table 2, and the propagated orbit is shown in Figure 3. The truth orbit was propagated using a lunar gravity model (up to degree and order 50), including third-body perturbations from the Earth and the Sun, as well as solar radiation pressure effects in a cannonball model with parameters $C_{R}=1.8$, satellite mass 850 kg, and 1.0 m2 cross-sectional area, following [LCRNS2025]. The simulation duration was set to 180 hours, correspocing to approximately 6 orbits.
@@
Figure 3: The simulated LCRNS (LDN-1) satellite orbit in Moon-centered inertial frame over 6 orbits (180 hours). The direction to Earth and Sun at the initial epoch is shown as blue and red arrows, respectively.
@@ 5.2 GNSS Orbits and Ephemerides
We considered a multi-constellation set of GPS, Galileo, and QZSS satellites. Truth GNSS orbits were generated using the IGS final orbit products [Johnston2017]. Specifically, SP3 precise ephemeris files provide satellite center-of-mass positions, which were interpolated using cubic splines to obtain satellite positions at the desired epochs. Antenna phase center information from ANTEX files was then used to map center-of-mass positions to antenna reference point (ARP) positions.
To emulate real-time navigation message performance, broadcast ephemerides were retrieved from daily multi-GNSS broadcast ephemeris files (via CDDIS). The broadcast ephemerides were used within the measurement models (Section 2) to introduce realistic ephemeris and clock modeling errors relative to the SP3 truth. The resulting GNSS constellation geometry at the simulation epoch is shown in Figure 4.
@@ 5.3 Link Budget Analysis and Satellite Visibility
where $P_{tx}$ and $G_{tx}$ is the transmit power and antenna gain of the GNSS satellites, $\theta_{tx},\varphi_{tx}$ are the transmit antenna off-boresite angle and azimuth, $\theta_{rx}$ are the receiver antenna off-boresite angle and azimuth, $G_{rx}$ is the receive antenna gain, $L_{path}$ is the free-space path loss, $L_{pol}$ is the polarization loss, $R_{loss}$ is the system implementation loss, $k_{b}=1.38\times 10^{-23}$ is the Boltzmann constant, and $T_{sys}$ is the system noise temperature.
The transmitter power of the GNSS satellites $P_{tx}$ is summarized in Table 3, which is based on the values in [Konitzer2022LuGRE] (from the ACE experiment and Magnetospheric Multiscale (MMS) Mission results) for GPS satellites, and [Nakajima2024] for QZSS. We also add an additional 3 dBW to the EIRP values for the L5 signals. Note that the value is not given for the Galileo satellites, since the Publications Office of the European Union [Menzione2024] publishes the EIRP (Effective Isotropic Radiated Power) values instead of the transmit power and the antenna gain separately. For the GNSS antenna patterns $G_{tx}$, we used the parameters from the NASA antenna characterization experiment (ACE) study [donaldson2020characterization] for Block II-F, the original data by Lockheed Martin for Block IIR and IIR-M satellites [Marquis2015], the United States Coast Guard Navigation Center [Fischer2022] for Block-III, Publications Office of the European Union [Menzione2024] for Galileo satellites, and the Cabinet Office of the Government of Japan for QZSS [QZSS2023]. The receiver antenna gain $G_{rx}$ has a peak gain of 14 dBi and a half-power beamwidth of 12.2 degrees, taken from [Delepaut2020GNSSLunar]. The antenna patterns of the selected GNSS satellites and the GNSS receiver we assumed for the lunar users are shown in Figure 5.

### v2

@@ 2.4.2 Multi-Constellation Inter-System Biases (ISB)
The above measurement equations are expressed in ITRF coordinates and GPS Time (GPST). For Galileo, the broadcast clock correction is provided in Galileo System Time (GST) using the Galileo ephemeris parameters $\boldsymbol{\alpha}_{\mathrm{GAL}}$. We convert the Galileo transmitter clock bias to GPST via
@@ 2.4.3 Light-Time Equation
$\bar{t}_{tx,i}^{\mathrm{GPST}}=t_{k}^{\mathrm{GPST}}-\frac{\left\|\mathbf{r}_{k}^{\mathrm{ITRF}}-\bar{\mathbf{r}}_{tx,i}^{\mathrm{ITRF}}(\bar{t}_{tx,i}^{\mathrm{GPST}},\boldsymbol{\alpha})\right\|}{c},$ (20)
where $t_{k}^{\mathrm{GPST}}$ and $\mathbf{r}_{k}^{\mathrm{ITRF}}$ are obtained from the lunar time/frame quantities via the transformations in (10). The computation of $\bar{\mathbf{r}}_{tx,i}^{\mathrm{ITRF}}$ from broadcast ephemerides follows [IS-GPS-200N].
@@ 2.4.4 Ionospheric and Plasmaspheric Delays and Mitigation
GNSS signals propagating from Earth to lunar orbit experience dispersive delays through the ionosphere and plasmasphere. The dominant first-order term can be removed by forming an ionosphere-free (IF) combination of dual-frequency measurements. For example, using L1 and L5 signals, the IF pseudorange is [kaplan2017understanding]
@@
where $\tilde{I}_{i,\mathrm{IF}}^{P}$ denotes the residual delay after first-order cancellation (e.g., higher-order terms and bending-related contributions), and
@@ 3 Stochastic-Cloning UD Filtering and Smoothing
This section presents a numerically robust filtering and smoothing framework for processing TDCP measurements in low-observability lunar navigation scenarios. While stochastic-cloning EKF have previously been proposed to accommodate delayed-state measurements [Iiyama2024TDCPNavigation], we observe that their direct covariance-propagation formulations are susceptible to numerical instability, particularly when processing low-noise TDCP observables.
@@ 3.3.2 UD rank-one update
$\mathbf{f}=(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}_{k,m}^{\top},\qquad\mathbf{g}=\hat{\mathbf{D}}^{-}\mathbf{f}.$ (37)
@@
$\hat{\mathbf{U}}^{+}\hat{\mathbf{D}}^{+}\bigl(\hat{\mathbf{U}}^{+}\bigr)^{\top}$ $=\hat{\mathbf{U}}^{-}\left[\hat{\mathbf{D}}^{-}-\frac{1}{a}\mathbf{g}\mathbf{g}^{\top}\right]\bigl(\hat{\mathbf{U}}^{-}\bigr)^{\top}.$ (38)
@@ 1 $\mathbf{f}\leftarrow(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}^{\top}$;
2 $\mathbf{g}\leftarrow\hat{\mathbf{D}}^{-}\mathbf{f}$;
@@ 3.4 Reducing Computation Cost for the Measurement Update
We first process the $m_{1}$ delayed-state measurements. After these updates, we obtain UD factors $(\hat{\mathbf{U}}^{+},\hat{\mathbf{D}}^{+})$ for the augmented covariance as described in Section 3.3. We then process the remaining $m_{2}$ current-state-only measurements. For such measurements, $\mathbf{H}_{k,m}^{(2)}=\mathbf{0}$, so the auxiliary vectors in Algorithm 2 $\mathbf{f}$ and $\mathbf{g}$ inherit a block structure that does not require the bottom-right UD factor $\hat{\mathbf{U}}^{-}_{22}$:
$\mathbf{f}$ $=(\hat{\mathbf{U}}^{-})^{\top}\mathbf{H}_{k,m}^{\top}=\begin{bmatrix}(\hat{\mathbf{U}}^{-}_{11})^{\top}\,(\mathbf{H}^{(1)}_{k,m})^{\top}\\ (\hat{\mathbf{U}}^{-}_{12})^{\top}\,(\mathbf{H}^{(1)}_{k,m})^{\top}\end{bmatrix},\qquad\mathbf{g}=\hat{\mathbf{D}}^{-}\mathbf{f}=\begin{bmatrix}\hat{\mathbf{D}}^{-}_{11}(\hat{\mathbf{U}}^{-}_{11})^{\top}(\mathbf{H}^{(1)}_{k,m})^{\top}\\ \hat{\mathbf{D}}^{-}_{22}(\hat{\mathbf{U}}^{-}_{12})^{\top}(\mathbf{H}^{(1)}_{k,m})^{\top}\end{bmatrix}$, (46)
@@
Importantly, the covariance update equation (45) does not depend on $\hat{\mathbf{U}}^{+}_{22}$, and from (46), the matrix $\hat{\mathbf{U}}^{-}_{22}$ ($\hat{\mathbf{U}}^{+}_{22}$ from the previous measurement) is not involved in the computation of $\mathbf{f}$ and $\mathbf{g}$, which is used to compute $\hat{\mathbf{U}}^{+}_{11}$, $\hat{\mathbf{U}}^{+}_{12}$, $\hat{\mathbf{D}}^{+}_{11}$, and $\hat{\mathbf{D}}^{+}_{22}$ in Algorithm 2. In addition, the Kalman gain for the first half of the augmented state vector (i.e., the current state) can be computed without using the second half of the Kalman gain vector, which depends on $b_{j},(j=n+1,\ldots,2n)$ in Algorithm 2.
@@ 3.5 Fixed-Interval Smoothing
The relationship between the estimated states and covariances in the forward and backward passes of the SC smoother is illustrated in Figure 3.
Figure 3: Relationship between the estimated states and covariances in backward smoothing
@@ 4.1.1 Time Update
$\frac{\partial f^{\text{dyn}}}{\partial\mathbf{x}}=F_{k}$ $=\begin{bmatrix}\mathbf{0}_{3\times 3}&\mathbf{I}_{3\times 3}&\mathbf{0}_{3\times 1}&\mathbf{0}_{3\times 1}&\mathbf{0}_{3\times 1}&\mathbf{0}_{3\times 1}\\ \frac{\partial\mathbf{a}}{\partial\mathbf{r}}&\frac{\partial\mathbf{a}}{\partial\mathbf{v}}&\mathbf{0}_{3\times 1}&\mathbf{0}_{3\times 1}&\mathbf{0}_{3\times 1}&\frac{\partial\mathbf{a}}{\partial\gamma}\\ \mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 3}&0&1&0&0\\ \left(\frac{\partial(c\dot{\delta}t_{rel})}{\partial\mathbf{r}}\right)^{\top}&\left(\frac{\partial(c\dot{\delta}t_{rel})}{\partial\mathbf{v}}\right)^{\top}&0&0&0&0\\ \mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 3}&0&0&1&0\\ \mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 3}&0&0&0&0\\ \mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 3}&0&0&0&0\\ \end{bmatrix}$ (60)
@@
$Q_{pos,vel}=\begin{bmatrix}\frac{1}{3}q_{a}\Delta t^{3}\mathbf{I}_{3\times 3}&\frac{1}{2}q_{a}\Delta t^{2}\mathbf{I}_{3\times 3}\\ \frac{1}{2}q_{a}\Delta t^{2}\mathbf{I}_{3\times 3}&q_{a}\Delta t\mathbf{I}_{3\times 3}\end{bmatrix}$ (63)
@@ 4.1.2 Measurement Preprocessing
    Signal Acquisition and Tracking: Acquire and track the GNSS signals to extract raw pseudorange and carrier phase measurements. The detailed algorithms for weak GNSS signal acquisition and tracking are beyond the scope of this paper. Readers are referred to [Capuano2016StandaloneGPS, Simone2023LuGREReceiver] for more information.
@@
    where $R_{E}$ is the Earth’s mean radius, $\mathbf{r}_{tx}$ is the position of the transmitting GNSS satellite in ECEF coordinates, and $\mathbf{r}$ is the position of the lunar satellite receiver in the ECEF coordinates. Remove all the measurements with tangential altitudes below a certain threshold to mitigate the impact of ionospheric delays.
@@
    Compute the receiver noise of the pseudorange and carrier phase measurements based on the observed carrier-to-noise ratio C/N0 values.
@@
    Compute the predicted measurements $\hat{z}_{k,i}$ and their residuals $y_{k,i}$ using the current state estimate $\hat{\mathbf{x}}_{k|k-1}$:
@@
    This criterion discards the TDCP measurements whenever the residual exceeds a 3-sigma bound, or exceeds the wavelength of the signal $\lambda_{L}$ (=19 cm for L1 and 24 cm for L5). It also discards all measurements where the 3-sigma bound itself exceeds the wavelength, which indicates insufficient state knowledge and/or measurement noise to reliably detect cycle slips. For this work, we only discard the TDCP measurements with detected cycle slips, and do not attempt to repair them.
@@ 4.1.3 Measurement Update
$\mathbf{H}_{k,i}$ $=\begin{bmatrix}-\frac{(\mathbf{r}_{k}-\mathbf{r}_{tx,k,i})^{\top}}{\|\mathbf{r}_{k}-\mathbf{r}_{tx,k,i}\|}&\mathbf{0}_{1\times 3}&1&0&0&0\end{bmatrix}\in\mathbb{R}^{1\times 10}$ (69)
@@ 5.1 Lunar Orbit
We used the initial conditions provided by [LCRNS2025] to simulate LCRNS satellites (LDN-1). Initial epoch of the simulation was set to March 1, 2025, 12:00:00 UTC, since the precise orbits, ephemeris, and the parameters needed for the ionospheric and plasmaspheric delay simulations were not available for the date in the reference (March 1, 2027, 00:00:00 UTC). To convert the initial conditions, the initial position and velocity vectors of the LCRNS satellite given at [LCRNS2025] were first converted to orbital elements at OP-frame (Orbital Plane frame) defined in [Ely2005FrozenLunarOrbits], and then the same initial orbital elements were used to generate the position and velocity vectors at the new initial epoch (March 1, 2025, 12:00:00 UTC) in the OP-frame. Finally, the position and velocity vectors were transformed back to the LCRF for the orbit propagation. The initial orbital elements in the OP-frame are summarized in Table 2, and the propagated orbit is shown in Figure 4. The truth orbit was propagated using a lunar gravity model (up to degree and order 50), including third-body perturbations from the Earth and the Sun, as well as solar radiation pressure effects in a cannonball model with parameters $C_{R}=1.8$, satellite mass 850 kg, and 1.0 m2 cross-sectional area, following [LCRNS2025]. The simulation duration was set to 180 hours, corresponding to approximately 6 orbits.
@@
Figure 4: The simulated LCRNS (LDN-1) satellite orbit in LCRF (Moon-centered inertial frame) over 6 orbits (180 hours). The direction to Earth and Sun at the initial epoch is shown as blue and red arrows, respectively.
@@ 5.2 GNSS Orbits and Ephemerides
We considered a multi-constellation set of GPS, Galileo, and Quasi-Zenith Satellite System (QZSS) satellites. Truth GNSS orbits were generated using the International GNSS Service (IGS) final orbit products [Johnston2017]. Specifically, final orbit products provide satellite center-of-mass positions, which were interpolated using cubic splines to obtain satellite positions at the desired epochs. Antenna phase center information from ANTEX files was then used to map center-of-mass positions to antenna reference point (ARP) positions.
To emulate real-time navigation message performance, broadcast ephemerides were retrieved from daily multi-GNSS broadcast ephemeris files via Crustal Dynamics Data Information System (CDDIS). The broadcast ephemerides were used within the measurement models (Section 2) to introduce realistic ephemeris and clock modeling errors relative to the truth orbits. The resulting GNSS constellation geometry at the simulation epoch is shown in Figure 5.
@@ 5.3 Link Budget Analysis and Satellite Visibility
where $P_{tx}$ and $G_{tx}$ is the transmit power and antenna gain of the GNSS satellites, $\theta_{tx},\varphi_{tx}$ are the transmit antenna off-boresite angle and azimuth, $\theta_{rx}$ are the receiver antenna off-boresite angle and azimuth, $G_{rx}$ is the receiver antenna gain, $L_{path}$ is the free-space path loss, $L_{pol}$ is the polarization loss, $R_{loss}$ is the system implementation loss, $k_{b}=1.38\times 10^{-23}$ is the Boltzmann constant, and $T_{sys}$ is the system noise temperature.
The transmitter power of the GNSS satellites $P_{tx}$ is summarized in Table 3, which is based on the values in [Konitzer2022LuGRE] (from the ACE experiment and Magnetospheric Multiscale (MMS) Mission results) for GPS satellites, and [Nakajima2024] for QZSS. We also add an additional 3 dBW to the Effective Isotropic Radiated Power (EIRP) values for the L5 signals. Note that the value is not given for the Galileo satellites, since the Publications Office of the European Union [Menzione2024] publishes the EIRP values instead of the transmit power and the antenna gain separately. For the GNSS antenna patterns $G_{tx}$, we used the parameters from the NASA antenna characterization experiment (ACE) study [donaldson2020characterization] for Block II-F, the original data by Lockheed Martin for Block IIR and IIR-M satellites [Marquis2015], the United States Coast Guard Navigation Center [Fischer2022] for Block-III, Publications Office of the European Union [Menzione2024] for Galileo satellites, and the Cabinet Office of the Government of Japan for QZSS [QZSS2023]. The receiver antenna gain $G_{rx}$ has a peak gain of 14 dBi and a half-power beamwidth of 12.2 degrees, taken from [Delepaut2020GNSSLunar]. The antenna patterns of the selected GNSS satellites and the GNSS receiver we assumed for the lunar users are shown in Figure 6.